# Harnessing Scale and Physics: A Multi-Graph Neural Operator Framework for PDEs on Arbitrary Geometries

### v1

@@ Abstract.
Partial differential equations, neural operator, geometric leaning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Artificial intelligence
@@ 3.2.2. Partitioning of Point Sets
To efficiently manage the challenge of generating overlapping partitions within a point set, we define each partition as a neighborhood ball within Euclidean space, characterized by centroid location and scale. We utilize a farthest point sampling (FPS) algorithm (Qi et al., 2017) to select centroids in a manner that ensures even coverage across the entire point set. This method starts with an arbitrary initial point and iteratively selects subsequent points that are the farthest in metric distance from all previously selected points. This approach not only guarantees a superior coverage compared to random sampling but also takes into account the spatial distribution of the points, contrasting with methods like convolutional neural networks (CNNs), that process data agnostically of spatial distributions. Consequently, our graph construction strategy establishes receptive fields in a data-dependent manner, significantly enhancing the model’s capability to capture and process complex spatial relationships. Detailed description of the FPS algorithm is provided in Appendix A.1.
@@ 3.2.3. Local Sampling
where $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ are the position vector of nodes $i$ and $j$, respectively. This metric also allows for connecting nodes that belong to different domains.
An illustrative example of this local sampling process and the resultant graph structure is shown in Figure 2(b). This visualization underscores how local sampling concentrates on areas rich in detail, thereby substantially enhancing the model’s capacity to learn nuanced behaviors and interactions, which are crucial for solving complex PDEs. Detailed description of the local sampling algorithm is provided in Appendix A.2.
@@ 3.4.1. Graph Attention Mechanism
The Graph Attention (GA) then compute the new representation for node $i$ as a weighted average of transformed features of its neighbors, applying a nonlinearity $\sigma$:
@@ A.1. Farthest Point Sampling Algorithm
Algorithm 1 Farthest Point Sampling Algorithm
1:Input: Coordinates Set $C$ of shape $(N,D)$, where $N$ is the number of nodes and $D$ is the dimensionality and $n$, the number of nodes to sample.
2:Initialize: $minDistances[N]\leftarrow\{\infty\};$
3:Initialize: $sampledIndices[n]\leftarrow\{-1\};$
4:$current\leftarrow\text{randint}(0,N-1);$triangleright Randomly select the first
5:$sampledIndices[0]\leftarrow current;$
6:for $i\leftarrow 1$ to $n-1$ do
7:     for $j\leftarrow 0$ to $N-1$ do
8:         $distance\leftarrow$ computeDistance $(C[j],C[current]);$
9:         if $distance<minDistances[j]$ then
10:              $minDistances[j]\leftarrow distance;$
11:         end if
12:     end for
13:     $current\leftarrow\text{argmax}(minDistances);$
14:     $SampledIndices[i]\leftarrow current;$
15:end for
16:Output: $sampledIndices$
@@ A.2. Local Sampling Algorithm
Local sampling plays a pivotal role in constructing graphs that capture the immediate neighborhood dynamics around nodes, crucial for addressing complex local interactions in PDE-solving environments. This method uses a high-frequency indicator (refer to Eq.(4)) to selectively sample nodes that are rich in detail, thereby ensuring the inclusion of only the most informative nodes in each local graph.
The sampling procedure involves computing a feature map that highlights areas of high detail. Nodes are then selected based on their prominence in this feature map, and the local connectivity among these nodes is established via a nearest neighbor approach using their positions. This ensures that each node’s local environment is effectively represented, capturing significant local interactions that are essential for precise PDE solutions.
The Local Sampling Algorithm, outlined in Algorithm 2, describes the process of selecting nodes and constructing a local graph. The algorithm takes as input the node features and their coordinates, processes them in batches, and performs local sampling and graph construction for each batch independently.
The topK function selects the $n$ highest values from the feature map within each batch, ensuring that nodes with the highest detail are chosen. The K-nearest neighbors (knn) function establishes edges based on the proximity of these selected nodes, effectively capturing local structural information. This targeted sampling and graph construction are vital for adapting the learning process to complex data variations and local interactions, thereby enhancing the model’s ability to solve PDEs with high precision.
@@
1:Input:
2:$X$: Node Set of shape $(N,D_{x})$, where $N$ is the number of nodes and $D_{x}$ is the dimensionality of node features.
3:$C$: Coordinates Set of shape $(N,D_{pos})$, where $D_{pos}$ is the dimensionality of positions.
4:$batch$: Array of batch indices corresponding to each node.
5:$n$: Number of nodes to sample per batch.
6:$k$: Number of nearest neighbors to consider.
7:Initialize: $sampledIndices\leftarrow[];$
8:Initialize: $sampledEdgeIndex\leftarrow[];$
9:$featureMap\leftarrow$ computeFeatureMap$(X,C,batch);$
10:for $i\leftarrow 1$ to $batch$ do
11:     $mask\leftarrow(batch==i);$
12:     $sampledIndices[i]\leftarrow$ topK$(FeatureMap[mask],n);$
13:     $sampledEdgeIndex[i]\leftarrow$ knn$(C[mask],C[mask],k);$
14:end for
15:Output: $sampledIndices$, $sampledEdgeIndex$
@@ Appendix C Details for Benchmarks
Figure 5. Visualization of a random sample from the Poisson dataset, showing the unstructured mesh, the input source term $f$ and the corresponding solution $u$.

### v3

@@ Abstract.
Partial differential equations, neural operator, geometric learning
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709173isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Artificial intelligence
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.14542736.
@@ 3.2.2. Partitioning of Point Sets
To efficiently manage the challenge of generating overlapping partitions within a point set, we define each partition as a neighborhood ball within Euclidean space, characterized by centroid location and scale. We utilize a farthest point sampling (FPS) algorithm (Qi et al., 2017) to select centroids in a manner that ensures even coverage across the entire point set. This method starts with an arbitrary initial point and iteratively selects subsequent points that are the farthest in metric distance from all previously selected points. This approach not only guarantees superior coverage compared to random sampling but also takes into account the spatial distribution of the points, contrasting with methods like convolutional neural networks (CNNs) that process data agnostically of spatial distributions. Consequently, our graph construction strategy establishes receptive fields in a data-dependent manner, significantly enhancing the model’s capability to capture and process complex spatial relationships. A detailed description of the FPS algorithm is provided in Appendix A.1.
@@ 3.2.3. Local Sampling
where $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ are the position vectors of nodes $i$ and $j$, respectively. This metric also allows for connecting nodes that belong to different domains.
An illustrative example of this local sampling process and the resultant graph structure is shown in Figure 2(b). This visualization underscores how local sampling concentrates on areas rich in detail, thereby substantially enhancing the model’s capacity to learn nuanced behaviors and interactions, which are crucial for solving complex PDEs. A detailed description of the local sampling algorithm is provided in Appendix A.2.
@@ 3.4.1. Graph Attention Mechanism
The Graph Attention (GA) then computes the new representation for node $i$ as a weighted average of transformed features of its neighbors, applying a nonlinearity $\sigma$:
@@ 6. Conclusion and Future Work
Here’s a more concise and consistent version of the acknowledgements:
Acknowledgements.
Wei Wang was supported by the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007), and the Guangzhou Municipal Science and Technology Projects (Nos. 2023A03J0003, 2023A03J0013, and 2024A03J0621). Zhilu Lai was supported by the Guangzhou-HKUST(GZ) Joint Funding Grant (No. 2023A03J0105), the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007), and the Guangzhou Municipal Science and Technology Project (No. 2023A03J0011).
@@ A.1. Farthest Point Sampling Algorithm
Input : Coordinates Set $C$ of shape $(N,D)$, where $N$ is the number of nodes and $D$ is the dimensionality; $n$, the number of nodes to sample.
Output : $sampledIndices$
$minDistances\leftarrow\{\infty\}$;
$sampledIndices\leftarrow\{-1\}$;
$current\leftarrow\text{randint}(0,N-1)$ $sampledIndices[0]\leftarrow current$;
for $i\leftarrow 1$ to $n-1$ do
       for $j\leftarrow 0$ to $N-1$ do
             $distance\leftarrow\text{computeDistance}(C[j],C[current])$;
             if $distance<minDistances[j]$ then
                   $minDistances[j]\leftarrow distance$;
             end if
       end for
      $current\leftarrow\text{argmax}(minDistances)$;
       $sampledIndices[i]\leftarrow current$;
end for
return $sampledIndices$;
Algorithm 1 Farthest Point Sampling Algorithm
@@ A.2. Local Sampling Algorithm
Local sampling is essential for constructing graphs that accurately capture the immediate neighborhood dynamics of nodes, which is vital for handling complex local interactions in PDE-solving environments. This method utilizes a high-frequency indicator (see Eq.(4)) to selectively sample detailed nodes, ensuring that each local graph includes only the most informative nodes. The sampling process begins by generating a feature map that highlights regions with high detail. Using the topK function, the top $n$ nodes with the highest feature values within each batch are selected. Subsequently, these selected nodes are connected using a K-nearest neighbors (knn) approach based on their spatial positions, effectively capturing the local structural information.
Algorithm 2 outlines the Local Sampling Algorithm, which operates as follows: given node features $X$, coordinates $C$, batch indices, the number of nodes to sample $n$, and the number of nearest neighbors $k$, the algorithm processes each batch by computing the feature map, selecting the top $n$ nodes using the topK function, and constructing edges between these nodes using the knn method. This targeted sampling and graph construction enable the model to adapt to complex data variations and local interactions, thereby enhancing its ability to solve PDEs with high precision.
Input :  $X$: Node Set of shape $(N,D_{x})$, where $N$ is the number of nodes and $D_{x}$ is the dimensionality of node features.;
$C$: Coordinates Set of shape $(N,D_{pos})$, where $D_{pos}$ is the dimensionality of positions.;
$batch$: Array of batch indices corresponding to each node.;
$n$: Number of nodes to sample per batch.;
$k$: Number of nearest neighbors to consider.
Output : $sampledIndices$, $sampledEdgeIndex$
$sampledIndices\leftarrow\emptyset$;
$sampledEdgeIndex\leftarrow\emptyset$;
$featureMap\leftarrow\text{computeFeatureMap}(X,C,batch)$;
for $i\leftarrow 1$ to $|\text{unique}(batch)|$ do
       $mask\leftarrow(batch=i)$;
      $sampledIndices[i]\leftarrow\text{topK}(featureMap[mask],n)$;
      $sampledEdgeIndex[i]\leftarrow\text{knn}(C[mask],C[mask],k)$;
end for
return $sampledIndices$, $sampledEdgeIndex$;
@@ Appendix C Details for Benchmarks
Figure 5. Visualization of a random sample from the Poisson dataset.